OpenStudy (anonymous):

Use the de Broglie relationship to determine the wavelength of a lithium atom moving at 2.6×10^5 m/s

OpenStudy (anonymous):

h According to the first de Broglie equation : wavelength = ------ , where m * v h is Planck's constant = 6.626 x 10^-34 (J s) m is the mass in kg v is velocity in m/s the mass of an atom = atomic mass / Avogadro's number for lithium it is = 6.941/ 6.022*10^23 = 1.153*10 ^ -23 g = 1.153* 10 ^-26 kg 6.626 x 10^-34 6.626 10^-34 wavelength = ----------------------- = ---------------- * ------------- (1.153*10^-26 * 2.6*10^5) 1.153*2.6 10^-26*10^5 6.626 wavelength = ------- * 10^-13 = 2.21 * 10 ^-13 m 2.998